OpenStudy (anonymous):

Find the distance between the points given. (2, 2) and (5, 5) 3√2 7√2

OpenStudy (anonymous):

7 sqrt 2?

OpenStudy (anonymous):

In order to find it, you should use the distance formula

OpenStudy (dtan5457):

Plug them in

OpenStudy (anonymous):

\[\sqrt{(x1-x2)^{2}+(y1-y2)^{2}}\]

OpenStudy (anonymous):

r the x the 5's?

OpenStudy (dtan5457):

\[\sqrt{(5-2)^2+(5-2)^2}=\sqrt{18}\]

OpenStudy (dtan5457):

How is 7 times the square root of 2, the square root of 18?

OpenStudy (dtan5457):

That would be the square root of 98

OpenStudy (dtan5457):

3√2 square the 3, multiply by the number in the radical-----radical 18

OpenStudy (anonymous):

oh k thnx